OpenStudy (anonymous):

PLEASE HELP! At a corner gas station, the revenue R varies directly with the number g of gallons of gasoline sold. If the revenue is $24.25 when the number of gallons sold is 5, find a linear equation that relates revenue R to the number of g of gallons of gasoline. then find the revenue R when the number of gallons of gasoline sold is 13.5.

OpenStudy (anonymous):

Hello, do you understand how to find the equation of a line?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

so how will the variables fit into the equation?

OpenStudy (anonymous):

Okay, we need two points to find the slope of the linear equation

OpenStudy (anonymous):

Our first point shall be (5, 24.25)

OpenStudy (anonymous):

Can you think of another point that we can intuitively know given the question?

OpenStudy (anonymous):

(x, 13.5) doesn't give the x value

OpenStudy (anonymous):

Okay, we know how much revenue the company makes when they sell 0 gas, correct?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

How much revenue will they make when they sell 0 gas?

OpenStudy (anonymous):

wouldn't it just be 0

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

the equation for the slope of a line is \[m=\frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (anonymous):

We now have two points (0,0) and (5, 24.25)

OpenStudy (anonymous):

Do you see how if we plug in these two points, we can obtain the slope m?

OpenStudy (anonymous):

slope would be 4.85

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

We know that the equation for a line is y = mx + b

OpenStudy (anonymous):

and we have the value of m

OpenStudy (anonymous):

we also have the value for y and x

OpenStudy (anonymous):

which are (5, 24.25)

OpenStudy (anonymous):

If we plug in these values, m, x and y, then all we have to find is b

OpenStudy (anonymous):

would the equation be y= 4.85x+48.5?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

24.25 = (4.85)(5)+b

OpenStudy (anonymous):

24.25 = 24.25+b

OpenStudy (anonymous):

Do you see what b must equal?

OpenStudy (anonymous):

b=0

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

We now have the full equation

OpenStudy (anonymous):

equation will be y=4.85x+0

OpenStudy (anonymous):

You got it

OpenStudy (anonymous):

Now you just need to find y = 4.85(13.5)

OpenStudy (anonymous):

wouldn't the 13.5 be in place of y and not x since the point is (x, 13.5)

OpenStudy (anonymous):

no, the amount of gas is your x variable

OpenStudy (anonymous):

the profit is your y value

OpenStudy (anonymous):

so the answer would be 65.47

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (anonymous):

You're welcome, good work!

OpenStudy (anonymous):

Did you need any help with any other questions?